hhanan:

Suppose you make coffee and it starts off just too hot to drink, so you put it in the refrigerator temporarily to cool it off. If the original temperature is 190 degrees Fahrenheit and the refrigerator is 36 degrees Fahrenheit, what will the temperature of the coffee be after 3 minutes? (k = 0.12)

ScamTheMan:

For this question we will use Newton's Law of Cooling to derive the answer. The function is: T(t)=Tₒ+(Tᵢ−Tₒ)∗e^(kt) T = Tempature t = Time Tₒ - Tempature of Surroundings Tᵢ - Intial Tempature k - Our constant that's been given We have all our values in the question above, it's just time to plug them in. T(3)=36+(190+36)∗e^(−0.36) Solving through we get our final answer: T(3)=143.4 Meaning after 3 minutes in a 36 Fahrenheit fridge, the coffee cools down to 143.4 degrees Fahrenheit

ScamTheMan:

I actually answered the same exact question yesterday, here's the link for that one: https://questioncove.com/study#/updates/6010c1c0894a92c766e9c000 I think you asked it too

hhanan:

oh sorry but thanks

ScamTheMan:

No problemo man, happy to help